# SymmCD: Symmetry-Preserving Crystal Generation with Diffusion Models

### v1

@@ Abstract
Generating novel crystalline materials has the potential to lead to advancements in fields such as electronics, energy storage, and catalysis. The defining characteristic of crystals is their symmetry, which plays a central role in determining their physical properties. However, existing crystal generation methods either fail to generate materials that display the symmetries of real-world crystals, or simply replicate the symmetry information from examples in a database. To address this limitation, we propose SymmCD111 Our code is publicly available at https://github.com/sibasmarak/SymmCD, a novel diffusion-based generative model that explicitly incorporates crystallographic symmetry into the generative process. We decompose crystals into two components and learn their joint distribution through diffusion: 1) the asymmetric unit, the smallest subset of the crystal which can generate the whole crystal through symmetry transformations, and; 2) the symmetry transformations needed to be applied to each atom in the asymmetric unit. We also use a novel and interpretable representation for these transformations, enabling generalization across different crystallographic symmetry groups.
We showcase the competitive performance of SymmCD on a subset of the Materials Project, obtaining diverse and valid crystals with realistic symmetries and predicted properties.
@@ 1 Introduction
Crystals serve as the fundamental building blocks of many materials, including most metals, ceramics, and rocks. The discovery of new crystalline materials is expected to lead to diverse technological breakthroughs in fields ranging from energy storage to computing hardware (Miret et al., 2024). Generative models have the potential to greatly accelerate this process by proposing new candidates materials, and possibly conditioning on desired properties or compositions.
@@
Figure 1: Illustration of the SymmCD method. Left. Representation of the unit cell of a 2D crystal with $p4m$ symmetry where the asymmetric unit and the site-symmetries of the atoms are highlighted. Leveraging symmetry results in a much more compact, yet complete representation. Right. Diffusion and denoising on the different components of the representation. For site symmetries and atom types, discrete diffusion is used. For the coordinate and asymmetric unit continuous diffusion is used. The diffusion and denoising processes preserve the space-group symmetry.
In this work, we propose a novel approach for generative modeling of inorganic crystals that ensures any desired distribution of space groups. The idea is similar to that of creating a paper snowflake, where we fold the paper to create an unconstrained space, and after an unconstrained cutting of the paper in this space, its unfolding creates an object with desired symmetries. In the context of crystals, the unconstrained space is called the asymmetric unit, which is a maximal subset of the unit cell with no redundancy. In order to be able to unfold the asymmetric unit, we need to generate the site symmetry of each atom inside the unit, i.e. the symmetry transformations that fix the atoms in place. In our generative process, the atomic positions are made consistent with generated site symmetries, enabling the unfolding of asymmetric unit into a symmetric crystal; see Figure 1.
Crystals and their individual atoms have many different types of symmetries, and so we need to address the issue of data-fragmentation. By representing symmetry information using standard crystallographic notations, such as Hermann–Mauguin notation or Wyckoff position labels (Hahn et al., 1983), we are faced with many crystals and site symmetries that have a low frequency in the training data. To address this problem, we introduce a novel representation of symmetries as binary matrices, which enables information-sharing and generalization across both crystal and site symmetries.
The main contributions of this work are as follows: I) We demonstrate a novel approach to generating crystals through the unconstrained generation of asymmetric units, along with their symmetry information. II) We introduce a physically-motivated representation for crystallographic site symmetries that generalizes across space groups. (III) We experimentally evaluate our method, finding that it performs on par with previous methods in terms of generating stable structures, while offering significantly improved computational efficiency due to our representation. (IV) We perform an in-depth analysis of the symmetry and diversity of crystal structures generated by existing generative models.
@@ 2 Related Work
In addition to structure-based modeling, prior work has also generated full-atom crystal structures, in which all atoms of the three-dimensional structure are generated. A range of generation methods including variational autoencoders (Noh et al., 2019; Xie et al., 2022), GANs (Nouira et al., 2018; Kim et al., 2020), reinforcement learning (Govindarajan et al., 2023), diffusion models (Zeni et al., 2023; Yang et al., 2023; Jiao et al., 2023; Klipfel et al., 2024), flow-matching models (Miller et al., 2024), and active learning based discovery (Merchant et al., 2023) have been used. These follow similar works in 3-D molecule generation (Hoogeboom et al., 2022; Garcia Satorras et al., 2021; Xu et al., 2022), but extend them by incorporating crystal periodicity. In addition to full-atom crystal generation, prior work has also applied text-based methods to understand and generate crystals using language models (Gruver et al., 2024; Flam-Shepherd & Aspuru-Guzik, 2023; Alampara et al., 2024).
Other works have pointed out the importance of symmetry of the generated structures. DiffCSP++ (Jiao et al., 2024), does so by using predefined structural templates from the training data and learning atomic types and coordinates compatible with the templates. While this is an interesting solution, we show that predefining the templates in this way severely limits the diversity and novelty of the generated samples. CrystalGFN (AI4Science et al., 2023) incorporates constraints on the lattice parameters and composition based on space groups, but does not guarantee that the atomic positions respect the desired symmetry. Finally, the concurrent works CrystalFormer (Cao et al., 2024) and WyCryst (Zhu et al., 2024) generate symmetric crystals by predicting atom symmetries. However, they simply use the labels of Wyckoff positions to encode symmetries, which does not enable generalization across groups. The methods are therefore limited to generating from space groups that are common in the dataset. By contrast, our method generalizes across groups and can generate valid crystals even from groups that are rare in the dataset.
@@ 3 Background
In addition to the translational symmetry of the lattice, crystals typically have many other symmetries. Understanding these symmetries is fundamental in characterizing crystals and directly relates to many of the properties of these materials. The space group $G$ of a crystal is the group of all Euclidean transformations that leave the crystal invariant, i.e., that simply permutes atoms of the same type. As space groups are subgroups of the Euclidean group, their elements can be represented as $\left(\mathbf{O},\mathbf{t}\right)$, where $\mathbf{O}\in O\left(n\right)$ and $\mathbf{t}\in\mathbb{R}^{3}$, with action on $\mathbf{x}\in\mathbb{R}^{3}$ defined as $\left(\mathbf{O},\mathbf{t}\right)\mathbf{x}=\mathbf{O}\mathbf{x}+\mathbf{t}$. The operations that are part of a space group can be generally understood as belonging to different types: translations, rotations, inversions, reflections, screw axes (combinations of rotations and translations), and glide planes (combinations of mirroring and translation). Different combinations of these symmetry operations are possible.
Two space-group belong to the same type if all their operations can be mapped to each other by an orientation-preserving Euclidean transformation (coordinate change). We denote the set of all space group types as $\mathcal{G}$. In 3 dimensions, there are only 230 unique space group types. By choosing a canonical coordinate system, we can in general work only with space group types.
The point group $P$ of a space group $G$ is the image of the homomorphism $\left(\mathbf{O},\mathbf{t}\right)\mapsto\mathbf{O}$, i.e the group obtained by keeping only the orthogonal parts of $G$. By contrast with space groups, any point group must at least preserve a single point, that is the origin. By a similar procedure to space groups, we can classify point groups and find that there are 32 crystallographic point groups types, consisting of inversions, rotations, and reflections. We denote the set of all point group types as $\mathcal{P}$.
@@
Points in the same orbit have conjugate site-symmetry groups. Therefore, site-symmetry groups related by conjugation can be understood as equivalent. This motivates a coarser level of classification that will be very useful. Two points $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{3}$ are part of the same Wyckoff position if their site-symmetry group is conjugate.
@@ 4.1 Representation of crystals with Wyckoff positions
As explained in the previous section, a crystal structure can, in general, be represented by the tuple ${\mathcal{C}}=\left(\mathbf{L},\mathbf{X},\mathbf{A}\right)$. This representation has been used in previous generative models for crystals (Xie et al., 2022; Jiao et al., 2023; Luo et al., 2023; Zeni et al., 2023). However, a fundamental limitation of a model based on this representation is that it does not leverage the inductive bias of crystal symmetry and offers no guarantees for the generated positions $\mathbf{X}$ and lattice $\mathbf{L}$ to satisfy anything but a trivial space group.
We introduce an alternative representation that respects symmetry in addition to having many desirable properties. First, we explicitly specify the space group type of the crystal $G\in\mathcal{G}$ in the representation. Given the space group, instead of representing each of the $N$ atoms individually with $\mathbf{X}\in\mathbb{R}^{3\times N}$ and $\mathbf{A}\in\mathbb{R}^{Z\times N}$, we represent the $M$ crystallographic orbits; replicating the atoms within the orbit then creates the crystal. As explained in Section 3, the Wyckoff position identifies a set of orbits by site-symmetries. Therefore, specifying the site-symmetry and an arbitrary orbit representative is sufficient to identify a crystallographic orbit. This corresponds to a representation of an asymmetric unit within the unit cell. We thus define the set of orbit representatives with their Wyckoff positions as the tuple ${\mathcal{C}}^{\prime}=\left({\mathbf{k}},\mathbf{X}^{\prime},{\mathbf{S}},\mathbf{A}^{\prime}\right)$, where ${\mathbf{k}}$ is a parametrization of the lattice (to be explained later), $\mathbf{X}^{\prime}=\left[\mathbf{x}^{\prime}_{1},\dots,\mathbf{x}^{\prime}_{M}\right]\in\mathbb{R}^{3\times M}$ are the representative’s fractional coordinates in the asymmetric unit, ${\mathbf{S}}=\left[S_{\mathbf{x}^{\prime}_{1}},\dots,S_{\mathbf{x}^{\prime}_{M}}\right]\in\mathcal{P}^{M}$ are the site-symmetry groups and $\mathbf{A}^{\prime}=\left[\mathbf{a}^{\prime}_{1},\dots,\mathbf{a}^{\prime}_{M}\right]\in\mathbb{R}^{Z\times M}$ are the atomic types.
@@
Our representation of crystals that explicitly takes into account symmetry is therefore given by the tuple $\mathcal{C}^{\prime}=\left(G,\mathbf{k},\mathbf{X}^{\prime},\mathbf{S},\mathbf{A}^{\prime}\right)$. We convert crystal structures to this representation using the spglib symmetry finding algorithm (Togo & Tanaka, 2018b) provided in the Pymatgen Python package (Ong et al., 2013).
@@ 4.2 Symmetry Representation
A key component of our representation of crystals with Wyckoff positions is the encoding of the space group $G$ and site-symmetry groups $\mathbf{S}^{\prime}$. While there are many existing methods to encode these symmetries, they generally do not make explicit the commonalities between the site-symmetries of Wyckoff positions in the same space group, and the commonalities between different space groups across crystal systems. This is an important limitation: because there are 230 space groups, not having a representation that is common across space groups results in dividing the effective amount of data the model is trained on by a large amount. For example, in the MP-20 dataset (Jain et al., 2013; Xie & Grossman, 2018) 113 space groups out of 169 in the training set have fewer than a hundred samples associated with them, and specific Wyckoff positions in each group have even fewer samples. We propose a method to represent the site-symmetries of different Wyckoff positions and to encode the symmetries of space groups to address this shortcoming.
We represent atom site-symmetries using a binary representation based on the oriented site-symmetry symbol used by the International Tables for Crystallography to describe Wyckoff positions (Hahn et al., 1983; Donnay & Turrell, 1974). The oriented site-symmetry symbols denote generators of the site-symmetry group along different possible axes, illustrated in Figure 2. In total, there are 15 possible axes of symmetry in a crystal, corresponding to each of the Cartesian axes, along with body and face diagonals.
Figure 3: SymmCD training and sampling pipeline. For training (green), the crystal structures are pre-processed to find the space group $G$ along with the site-symmetries $\mathbf{S}$ and a set of orbit representatives inside the asymmetric unit. The denoising model is a GNN with fully connected graphs, followed by a decoder. For sampling (blue), the positions are projected to the closest one compatible with their site-symmetry. Then, the asymmetric unit is replicated to obtain the unit cell.
@@
The space group $G$ can also be encoded into a binary representation using a similar scheme, by listing out the 15 possible axes of symmetry and listing out the possible symmetry operations along each axis. Unlike the point group symmetries of atoms, these space group symmetry operations may involve translations and so include screw and glide transformations, leading to 26 possible symmetry operations. Further details are included in Appendix C. As a part of this project, both the site-symmetry representations and the space group representations have been incorporated as functions in the PyXtal software library (Fredericks et al., 2021).
@@ 4.3 Diffusion Model
We consider a separate diffusion process over the different components of the crystal representation. We apply discrete diffusion from Austin et al. (2021) for site-symmetries and atom types. Rather than adding Gaussian noise as in conventional diffusion, we add noise to categorical features by multiplying probability vectors by a transition matrix and sampling from the new probabilities. Inspired by Vignac et al. (2023), the transition matrices are parameterized so that the process converges to the marginals from the data distribution for atom types and site-symmetries. The loss function used for discrete diffusion on atomic types is
@@
where ${\mathbf{a}}_{0}$ is the initial one-hot encoding of the atom types for a single representative and $\bar{\mathbf{Q}}_{t}=\prod_{i=1}^{t}\mathbf{Q}_{i}\in\mathbb{R}^{Z\times Z}$ is the cumulative product of transition matrices between timesteps, and $\hat{{\mathbf{a}}}_{i}$ are the predicted denoised probabilities. The same loss function is used for site-symmetries.
@@
$\mathcal{L}_{\mathbf{k}}=\mathbb{E}_{\epsilon_{{\mathbf{k}}}\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathrm{U}(1,T)}[||m\odot\epsilon_{{\mathbf{k}}}-\hat{\epsilon}_{{\mathbf{k}}}({\mathcal{C}}^{\prime}_{t},t)||_{2}^{2}],$
@@
The architecture of the denoiser is a message-passing graph neural network that operates on a fully connected graph of representatives, based on Jiao et al. (2023). Features for each representative ${\mathbf{h}}_{i}$ are initialized using an embedding of their atom types ${\mathbf{a}}_{i}$ and their site symmetries ${\mathbf{S}}_{i}$, along with the graph-level features of the diffusion timestep $t$, the lattice features ${\mathbf{k}}$, and an embedding of the space group $G$. At each layer, messages ${\mathbf{m}}_{ij}$ are computed between representatives $i$ and $j$ by applying an MLP to ${\mathbf{h}}_{i},{\mathbf{h}}_{j}$, and a Fourier basis embedding of the vector ${\mathbf{x}}_{i}-{\mathbf{x}}_{j}$ to respect periodic invariance. These messages are then used to update ${\mathbf{h}}_{i}$. More details on the architecture are included in Section E.1. Note that this denoising network is not equivariant. It is not necessary since the unit cell axes provide a canonical reference system (Kaba et al., 2023). We also found that using an equivariant denoising network like E$(n)$-GNN did not work well in part due to the fact that since we use periodic encodings, the crystal structure input has a translational symmetry. An equivariant model is not be able to break that symmetry (Kaba & Ravanbakhsh, 2023) resulting in a inability to output correct positions in the asymmetric unit (or unit cell).
@@
Algorithm 1 Training the Crystal Generation Diffusion Model
@@
Algorithm 2 Sampling from Crystal Generation Diffusion Model

### v2

@@ Abstract
Generating novel crystalline materials has the potential to lead to advancements in fields such as electronics, energy storage, and catalysis. The defining characteristic of crystals is their symmetry, which plays a central role in determining their physical properties. However, existing crystal generation methods either fail to generate materials that display the symmetries of real-world crystals, or simply replicate the symmetry information from examples in a database. To address this limitation, we propose SymmCD111 Our code is publicly available at https://github.com/sibasmarak/SymmCD/, a novel diffusion-based generative model that explicitly incorporates crystallographic symmetry into the generative process. We decompose crystals into two components and learn their joint distribution through diffusion: 1) the asymmetric unit, the smallest subset of the crystal which can generate the whole crystal through symmetry transformations, and; 2) the symmetry transformations needed to be applied to each atom in the asymmetric unit. We also use a novel and interpretable representation for these transformations, enabling generalization across different crystallographic symmetry groups. We showcase the competitive performance of SymmCD on a subset of the Materials Project, obtaining diverse and valid crystals with realistic symmetries and predicted properties.
@@ 1 Introduction
Crystals serve as the fundamental building blocks of many materials, and the discovery of new crystalline materials is expected to lead to diverse technological breakthroughs in fields ranging from energy storage to computing hardware (Miret et al., 2024). Generative models have the potential to greatly accelerate this process by proposing new candidates materials, and possibly conditioning on desired properties or compositions.
@@
Figure 1: Illustration of the SymmCD method. Left. Representation of the unit cell of a 2D crystal with $p4m$ symmetry where the asymmetric unit and the site symmetries of the atoms are highlighted. Leveraging symmetry results in a much more compact, yet complete representation. Right. Diffusion and denoising on the different components of the representation. For site symmetries and atom types, discrete diffusion is used. For the coordinate and asymmetric unit continuous diffusion is used. The diffusion and denoising processes preserve the space group symmetry.
In this work, we propose a novel approach for generative modeling of inorganic crystals that ensures any desired distribution of space groups. The idea is similar to that of creating a paper snowflake, where we fold the paper to create an unconstrained space, and after an unconstrained cutting of the paper in this space, its unfolding creates an object with desired symmetries. In the context of crystals, the unconstrained space is called the asymmetric unit, which is a maximal subset of the unit cell with no redundancy. To unfold the asymmetric unit, we need to generate the site symmetry of each atom inside the unit, i.e. the symmetry transformations that fix the atoms in place. In our generative process, the atomic positions are made consistent with generated site symmetries, enabling the unfolding of an asymmetric unit into a symmetric crystal; see Figure 1.
Crystals and their individual atoms have many different types of symmetries, so we need to address the issue of data-fragmentation. By representing symmetry information using standard crystallographic notations, such as Hermann–Mauguin notation or Wyckoff position labels (Hahn et al., 1983), we are faced with many crystals and site symmetries that have a low frequency in the training data. To address this problem, we introduce a novel representation of symmetries as binary matrices, which enables information-sharing and generalization across both crystal and site symmetries.
The main contributions of this work are as follows: I) We demonstrate a novel approach to generating crystals through the unconstrained generation of asymmetric units, along with their symmetry information. II) We introduce a physically-motivated representation for crystallographic site symmetries that generalizes across space groups. III) We experimentally evaluate our method, finding that it performs on par with previous methods at generating stable structures, while offering significantly improved computational efficiency due to our representation. IV) We analyse the symmetry and diversity of crystal structures generated by existing generative models.
@@ 2 Related Work
In addition to structure-based modeling, prior work has also generated full-atom crystal structures, in which all atoms of the three-dimensional structure are generated. A range of generation methods including VAEs (Noh et al., 2019; Xie et al., 2022), GANs (Nouira et al., 2018; Kim et al., 2020), reinforcement learning (Govindarajan et al., 2023), diffusion models (Zeni et al., 2023; Yang et al., 2023; Jiao et al., 2023; Klipfel et al., 2024), flow-matching models (Miller et al., 2024), and active learning based discovery (Merchant et al., 2023) have been used. These follow similar works in 3D molecule generation (Hoogeboom et al., 2022; Garcia Satorras et al., 2021; Xu et al., 2022), but extend them by incorporating crystal periodicity. In addition to full-atom crystal generation, prior work has also applied text-based methods to understand and generate crystals using language models (Flam-Shepherd & Aspuru-Guzik, 2023; Gruver et al., 2024; Alampara et al., 2024).
Other works have pointed out the importance of symmetry of the generated structures. DiffCSP++ (Jiao et al., 2024), does so by using predefined structural templates from the training data and learning atomic types and coordinates compatible with the templates. While this is an interesting solution, we show that predefining the templates in this way severely limits the diversity and novelty of the generated samples. CrystalGFN (AI4Science et al., 2023) incorporates constraints on the lattice parameters and composition based on space groups but does not guarantee that the atomic positions respect the desired symmetry. Finally, the concurrent works CrystalFormer (Cao et al., 2024) and WyCryst (Zhu et al., 2024) generate symmetric crystals by predicting atom symmetries. However, they directly use the labels of Wyckoff positions to encode symmetries, which does not enable generalization across groups. The methods are therefore limited to generating from space groups that are common in the dataset. By contrast, our method generalizes across groups and can generate valid crystals even from groups that are rare in the dataset.
@@ 3 Background
In addition to the translational symmetry of the lattice, crystals typically have many other symmetries. Understanding these symmetries is fundamental in characterizing crystals and directly relates to many of the properties of these materials. The space group $G$ of a crystal is the group of all Euclidean transformations that leave the crystal invariant, i.e., that simply permutes atoms of the same type. As space groups are subgroups of the Euclidean group, their elements can be represented as $\left(\mathbf{O},\mathbf{t}\right)$, where $\mathbf{O}\in O\left(n\right)$ and $\mathbf{t}\in\mathbb{R}^{3}$, with action on $\mathbf{x}\in\mathbb{R}^{3}$ defined as $\left(\mathbf{O},\mathbf{t}\right)\mathbf{x}=\mathbf{O}\mathbf{x}+\mathbf{t}$. The operations that are part of a space group can be generally understood as belonging to different types: translations, rotations, inversions, reflections, screw axes (combinations of rotations and translations), and glide planes (combinations of mirroring and translation).
Two space groups belong to the same type if all their operations can be mapped to each other by an orientation-preserving Euclidean transformation (coordinate change). We denote the set of all space group types as $\mathcal{G}$. In 3 dimensions, there are only 230 unique space group types. By choosing a canonical coordinate system, we can in general work only with space group types. The point group $P$ of a space group $G$ is the image of the homomorphism $\left(\mathbf{O},\mathbf{t}\right)\mapsto\mathbf{O}$, i.e the group obtained by keeping only the orthogonal parts of $G$. By contrast with space groups, any point group must at least preserve a single point, that is the origin. By a similar procedure to space groups, we can classify point groups and find that there are 32 crystallographic point group types, consisting of inversions, rotations, and reflections. We denote the set of all point group types as $\mathcal{P}$.
@@ 4.1 Representation of crystals with Wyckoff positions
As explained in the previous section, a crystal structure can in general be represented by the tuple ${\mathcal{C}}=\left(\mathbf{L},\mathbf{X},\mathbf{A}\right)$. This representation has been used in previous generative models for crystals (Xie et al., 2022; Jiao et al., 2023; Luo et al., 2023; Zeni et al., 2023). However, a fundamental limitation of a model based on this representation is that it does not leverage the inductive bias of crystal symmetry and offers no guarantees for the crystal to satisfy anything but a trivial space group.
We introduce an alternative representation that respects symmetry in addition to having many desirable properties. First, we explicitly specify the space group type of the crystal $G\in\mathcal{G}$ in the representation. Given the space group, instead of representing each of the $N$ atoms individually with $\mathbf{X}\in\mathbb{R}^{3\times N}$ and $\mathbf{A}\in\mathbb{R}^{Z\times N}$, we represent the $M$ crystallographic orbits; replicating the atoms within the orbit then creates the crystal. As explained in Section 3, the Wyckoff position identifies a set of orbits by site symmetries. Therefore, specifying the site symmetry and an arbitrary orbit representative is sufficient to identify a crystallographic orbit. This corresponds to a representation of an asymmetric unit within the unit cell. We thus define the set of orbit representatives with their Wyckoff positions as the tuple ${\mathcal{C}}^{\prime}=\left({\mathbf{k}},\mathbf{X}^{\prime},{\mathbf{S}},\mathbf{A}^{\prime}\right)$, where ${\mathbf{k}}$ is a parametrization of the lattice (to be explained later), $\mathbf{X}^{\prime}=\left[\mathbf{x}^{\prime}_{1},\dots,\mathbf{x}^{\prime}_{M}\right]\in\mathbb{R}^{3\times M}$ are the representative’s fractional coordinates in the asymmetric unit, ${\mathbf{S}}=\left[S_{\mathbf{x}^{\prime}_{1}},\dots,S_{\mathbf{x}^{\prime}_{M}}\right]\in\mathcal{P}^{M}$ are the site symmetry groups and $\mathbf{A}^{\prime}=\left[\mathbf{a}^{\prime}_{1},\dots,\mathbf{a}^{\prime}_{M}\right]\in\mathbb{R}^{Z\times M}$ are the atomic types.
@@
Our representation of crystals that explicitly takes into account symmetry is therefore given by the tuple $\mathcal{C}^{\prime}=\left(G,\mathbf{k},\mathbf{X}^{\prime},\mathbf{S},\mathbf{A}^{\prime}\right)$. We convert crystal structures to this representation using the spglib symmetry finding algorithm (Togo & Tanaka, 2018b) provided by Pymatgen (Ong et al., 2013).
@@ 4.2 Symmetry Representation
A key component of our representation of crystals with Wyckoff positions is the encoding of the space group $G$ and site symmetry groups $\mathbf{S}^{\prime}$. While there are many existing methods to encode these symmetries, they generally do not make explicit the commonalities between the site symmetries of Wyckoff positions in the same space group, and the commonalities between different space groups across crystal systems. This is an important limitation: because there are 230 space groups, not having a representation that is common across space groups results in dividing the effective amount of data the model is trained on by a large amount. For example, in the MP-20 dataset (Jain et al., 2013; Xie & Grossman, 2018) 113 space groups out of 169 in the training set have fewer than a hundred samples associated with them, and specific Wyckoff positions in each group have even fewer samples. We propose a method to represent the site symmetries of different Wyckoff positions and to encode the symmetries of space groups to address this shortcoming.
We represent atom site symmetries using a binary representation based on the oriented site symmetry symbol used by the International Tables for Crystallography to describe Wyckoff positions (Hahn et al., 1983; Donnay & Turrell, 1974). The oriented site symmetry symbols denote generators of the site symmetry group along different possible axes, illustrated in Figure 2. In total, there are 15 possible axes of symmetry in a crystal, corresponding to each of the Cartesian axes, along with body and face diagonals.
@@
The space group $G$ can also be encoded into a binary representation using a similar scheme, by listing out the 15 possible axes of symmetry and listing out the possible symmetry operations along each axis. Unlike the point group symmetries of atoms, these space group symmetry operations may involve translations and so include screw and glide transformations, leading to 26 possible symmetry operations. Further details are included in Appendix 4.1. As a part of this project, both the site symmetry representations and the space group representations have been incorporated as functions in the PyXtal software library (Fredericks et al., 2021).
@@ 4.3 Diffusion Model
Figure 3: SymmCD training and sampling pipeline. For training (green), the crystal structures are pre-processed to find the space group $G$ along with the site symmetries $\mathbf{S}$ and a set of orbit representatives inside the asymmetric unit. The denoising model is a GNN with fully connected graphs, followed by a decoder. For sampling (blue), the positions are projected to the closest one compatible with their site symmetry. Then, the asymmetric unit is replicated to obtain the unit cell.
@@
We consider a separate diffusion process over the different components of the crystal representation. We apply discrete diffusion from Austin et al. (2021) for site symmetries and atom types. Rather than adding Gaussian noise as in conventional diffusion, we add noise to categorical features by multiplying probability vectors by a transition matrix and sampling from the new probabilities. Inspired by Vignac et al. (2023), the transition matrices are parameterized so that the process converges to the marginals from the data distribution for atom types and site symmetries. The loss function used for discrete diffusion on atomic types is
@@
where ${\mathbf{a}}_{0}$ is the initial one-hot encoding of the atom types for a single representative and $\bar{\mathbf{Q}}_{t}=\prod_{i=1}^{t}\mathbf{Q}_{i}\in\mathbb{R}^{Z\times Z}$ is the cumulative product of transition matrices between timesteps, and $\hat{{\mathbf{a}}}_{i}$ are the predicted denoised probabilities. The same loss function is used for site symmetries.
@@
$\mathcal{L}_{\mathbf{k}}=\mathbb{E}_{\epsilon_{{\mathbf{k}}}\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathrm{U}(1,T)}\left[||m\odot\epsilon_{{\mathbf{k}}}-\hat{\epsilon}_{{\mathbf{k}}}({\mathcal{C}}^{\prime}_{t},t)||_{2}^{2}\right],$
@@
The architecture of the denoiser is a message-passing neural network that operates on a fully connected graph of representatives, based on Jiao et al. (2023). Features for each representative ${\mathbf{h}}_{i}$ are initialized using an embedding of their atom types ${\mathbf{a}}_{i}$ and their site symmetries ${\mathbf{S}}_{i}$, along with the graph-level features of the diffusion timestep $t$, the lattice features ${\mathbf{k}}$, and an embedding of the space group $G$. At each layer, messages ${\mathbf{m}}_{ij}$ are computed between representatives $i$ and $j$ by applying an MLP to ${\mathbf{h}}_{i},{\mathbf{h}}_{j}$, and a Fourier embedding of the vector ${\mathbf{x}}_{i}-{\mathbf{x}}_{j}$ to respect periodic invariance. These messages are then used to update ${\mathbf{h}}_{i}$. More details on the architecture are included in Section E.1. Note that the denoising network is not equivariant. It is not necessary since the unit cell axes provide a canonical reference system (Kaba et al., 2023). We also found that using an equivariant denoising network like EGNN Satorras et al. (2021) did not work well in part due to the fact that since we use periodic encodings, the crystal structure input has a translational symmetry. An equivariant model may not be able to break that symmetry (Kaba & Ravanbakhsh, 2023) resulting in a inability to output correct positions in the asymmetric unit (or unit cell).
@@
Algorithm 1 Training SymmCD
@@
Algorithm 2 Sampling from SymmCD